3.- (10 points) Two parall 10 A. If wo parallel wires are 10.0 cm apart, and each carries a current of crents are in the same direction, find the force per unit of length oy one of the wires on the other. Are the wires attracted or repelled? 个 ocm
Answer #1

Forec per unit length of wire 1 on 2

\frac{\Delta F}{\Delta l}=\frac{\mu_{0}I_{1}I_{2}}{2\pi r}

\frac{\Delta F}{\Delta l}=\frac{(4\pi \times 10^{-7})(10)(10)}{2\pi (0.1)}=2\times 10^{-4}N/m

Since the current are in the same direction, two wires attract each other